OpenStudy (anonymous):

Is the equation true, false, or open? 5x+2=6x+11 WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (texaschic101):

5x + 2 = 6x + 11 5x - 6x = 11 - 2 -x = 9 x = -9 check... 5(-9) + 2 = 6(-9) + 11 -45 + 2 = - 54 + 11 -43 = -43 (correct) It is true